OpenStudy (anonymous):

what is the second term of (p+q)^5 1. 25p^4 2. 5p^4q 3. 25p^4q 4. 125p^4q

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thank you. can you help me with another ! I SUCKKKKKK AT ALGEBRA

OpenStudy (anonymous):

@Lolovitz

OpenStudy (anonymous):

No problem bro ( or sis actually ;] )

OpenStudy (anonymous):

thank you. its a major help :) One zero of f(x)=x^3+2x^2-5x-6 is 2. what re other zeroes of the function?

OpenStudy (anonymous):

1. 1 &3 2. -1 & -3 3. 2 & -3 4. 2 & 3

OpenStudy (anonymous):

If you use x=1 f(x) =1+2-5-6 which is not equal to zero. So 1 is not the zero of this function. Now we will see -3. x=3 f(x)=-27+18+15-6=0 So -3 is the zero of this function. So it is the thirs answer as the only one without 1 and with -3. 3.2 & -3 Hope i helped. Glad to help with anymore problematic question you may have.

OpenStudy (anonymous):

@Lolovitz OH MY GOSH!!!!!! you made it so much easier to understand.... ive been so confused. so I just plug the #"s into the x?

OpenStudy (anonymous):

Yeah. Once you substitue the X with real values you can calculate the value of function for this specific value of x. Needless to say if for a specific value of x f(x)=0 then this value of x is the zero of this function. However if you have a function and you have no possible idea's what the possible zero's of this function are you can take it to pieces in some cases. For exampl the function f(x)=x^3-2x^2-5x+6=(x-3)(x+2)(x-1) so the zero's of this function are the values of x in which x-3=0 OR x+2=0 OR x-1=0 So x=3 OR x=-2 OR x=1 Hope i helped. With any more question just message me somehow ;] :3

OpenStudy (anonymous):

thank you :) I have one more question from my homework, it asks which of the 2 x's fits as its only 2 real zeroes, but I don't understand how you put two into one equation ! which f the following quartic functions has x=-2 and x=-3 as its only two real zeroes?

OpenStudy (anonymous):

well If x=-2 and x=-3 are the zero's then x+2=0 or x+3=0 f(x)=(x+3)(x+2)*SUMTHING=(x^2+5x+6)*SUMTHING

OpenStudy (anonymous):

1. x^4-5x^3+7x^2-5x+6 2. X^4+5x^3 +7x^2+5x+6 3. x^4+5x^3+7x^2-5x-6 4. x^4-5x^3+7x^2+5x-6

OpenStudy (anonymous):

hmm. Just plug -3 and -2 seperatly into each one of this function and see which one for x=-2 equals zero and for x=-3 equals zero.

OpenStudy (anonymous):

so both answers have to = 0 when I separately plug in -2 then -3?

OpenStudy (anonymous):

Yes. For any more question find me on skype if you can - marek.wojnicz the guy with killer bunny on the avatar.;]